OpenStudy (kitten_is_back):

A silver piece of art in the shape of a regular heptagon has an apothem length of 27 mm and a side length of 26 mm. What is its area? mm2

OpenStudy (welshfella):

we can split this up into seven equal triangles whose height is 27 and base 26 mm

OpenStudy (kitten_is_back):

@jabez177

jabez177 (jabez177):

What @welshfella was saying. So first multiply 27 x 26, what is your answer?

OpenStudy (welshfella):

the area of each triangle = 1/2 * base * height = 1/2 * 26 * 27

OpenStudy (welshfella):

after you have worked that out the area of the heptagon = 7 times that.

OpenStudy (kitten_is_back):

lagging

OpenStudy (welshfella):



OpenStudy (welshfella):

the apotherm 27 is also the height of each triangle

OpenStudy (welshfella):

the vertex of each triangle is also the center of the regular heptagon ( 7 - sides)

OpenStudy (kitten_is_back):

702

OpenStudy (welshfella):

understand?

OpenStudy (kitten_is_back):

yes

OpenStudy (welshfella):

1/2 * 26 * 27 * 7

OpenStudy (welshfella):

its not 702

OpenStudy (welshfella):

i think yiuve worked out 1/2 * 26*27 * 2 - there are 7 triangles altogether